Question 5. A train carrying 50,000 kg of methanol (CH OH) derails, spilling its cargo. (a) How many kg of oxygen would be required to burn this much methanol? (b) How much oxygen would be required if aerobic microorganisms are used to degrade the methanol (assume that fe 0.7)? nitrate must be provided (as kg of nitrate)? supplied to biodegrade the methanol (assume nitrate is the nitrogen source for cell synthesis and f,-0.7). (c) For aerobic biodegradation, how much cell mass will be produced if nitrate is the nitrogen source? How much (d) Suppose denitrifying organisms are used to degrade the methanol. How much nitrate (as kg of nitrate) must be
Answer #1

Q5 a)

Mass of Methanol=50,000 Kg=50000000g=(5/34)*10^7moles (Given mass/Molecular mass=Number of moles)(Molecular mass of Methanol=34)

CH3OH + (3/2) O2 \rightarrow (2)H2O+CO2

Therefore 1 mole od=f methanol will require 1.5 moles of Oxygen for complete combustion.

Moles of Oxygen required=(5/34)*10^7*1.5 moles

Mass of 1 mole of Oxygen=32

mass of Oxygen required=(5/34)*10^7*1.5*32=70588.235Kg